OpenStudy (sierra_grey):

(a)An angle measures 41 degrees. What is the measure of its complement? (b)An angle measures 35 degrees. What is the measure of its supplement?

OpenStudy (anonymous):

property of complementary angles is that its sum gives us 90degrees now here one angle is 41deg so the second is 90-41

OpenStudy (anonymous):

now you can guess the complement of 41 degrees

OpenStudy (sierra_grey):

Compliment is 49 and Supplement is 145?

OpenStudy (sierra_grey):

Complement*

OpenStudy (anonymous):

See compliemt of a given angles is given by the formula: Complement = (90 - Degrees) And. Suppliment (180 - Degress) SO a) (90 - 41) = 49[Complement] (180 - 41) = 139[Supplement] b.)(90 -35) = 55[Complement] (180 - 35) = 145[Supplement]